OpenStudy (anonymous):

Please help me understand this....-6= 4 - 2t + 4t + 16, what does t equal, what is the procedure step by step please, I have 16 more of these to do, I just need to learn how!

OpenStudy (anonymous):

-6=4-2t+4t+16 combine like terms: -6=2t+20 then subtract 20 from both sides -26=2t then divide both sides by 2 -13=t

OpenStudy (anonymous):

let me know if youre still confused =D

OpenStudy (anonymous):

OMGGGGG

OpenStudy (anonymous):

Thank you, thank you soo much! ill pm you if theres something I dont understand! thanks!

OpenStudy (anonymous):

no problem =D